OpenStudy (anonymous):

Simplify the Expression, if possible 2x^2/6x^3-10x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\frac{2x^2}{6x^3-10x}=\frac{2x}{6x^2-10}=\frac{2}{3x^2-5}\]